OpenStudy (anonymous):

What is 1/e eqivalent to?

Directrix (directrix):

It's e^(-1) but I don't think that's the answer you are seeking. As a series, see attachment.

OpenStudy (anonymous):

Yes, directrix, it's not what I'm looking for... but thanks for trying :)

OpenStudy (campbell_st):

e is approximately 2.73.... its a transcendental number like pi so e^(-1) = 1/2.73...

OpenStudy (campbell_st):

oops thats e^1

Directrix (directrix):

For what are you looking about 1/e?

OpenStudy (anonymous):

I'm seeing if it equals 0, so that it can pass through the origin. Eg.. y=1/e passes through the origin..

Directrix (directrix):

We are not talking y = 1/(e^x). Correct.

OpenStudy (campbell_st):

no.... its never zero.... because you graph e^x negative powers mean 1/e^x.... the only way the graph goes below the x-axis is for an equation such as y = e^x - b where b is a constant

OpenStudy (campbell_st):

pick up a calculator... it has a function for e

OpenStudy (anonymous):

yes, but when I do 1/e, it gives syntax error.

OpenStudy (campbell_st):

enter 1/e1

OpenStudy (anonymous):

Oh. Ok. 0.3679 it's giving me.

OpenStudy (campbell_st):

that 1/e^1 then 1/e2 will be smaller 1/e0 = 1

OpenStudy (anonymous):

yes, it's true

OpenStudy (campbell_st):

hope that helps

OpenStudy (anonymous):

Have you seen this http://www.wolframalpha.com/input/?i=y%3D1%2Fe#

OpenStudy (campbell_st):

if you have a graphing program of calculator... graph y = e^x and then y = e^(-x)

OpenStudy (anonymous):

Oh, that helps @FoolForMath I don't have a graphing program... except wolfram

OpenStudy (campbell_st):

just enter graph y = e^x etc

OpenStudy (anonymous):

@order: Apart from wolframalpha, Mathematica, Geogebra and Grapeq are good choices.

OpenStudy (anonymous):

Thanks! :-)

OpenStudy (anonymous):

Glad to help :)